The height of the students in a certain class is following normal distribution with mean height as 165 cm and standard deviation of 25 cm. There are 60 students in that class. Determine
Find the lowest height among the tallest 5 student in a class
1
Expert's answer
2020-06-22T18:16:33-0400

5 tallest students is 5/60=8.3% in a class

P(x>X)=1-P(x<X)=0.083

P(x<X)=1-0.083=0.917, using table, z-value for it equals to 1.39

(X-μ)/σ =1.39

(x-165)/25=1.39

x=165+25*1.39=199.75

So, the lowest height among the tallest 5 students in a class equals to 200cm